The number of internal disk drives​ (in millions) made at a plant in Taiwan during the past 5 years​ follows:

                                                                     

Year

Disk Drives

1

140140

2

156156

3

190190

4

202202

5

216216

​a) Using simple linear regression..., the forecast for the number of disk drives to be made next year​ = 240.2

disk drives ​(round your response to one decimal​ place).

​b) The mean squared error​ (MSE) when using simple linear regression​ = ????drives^2

​(round your response to one decimal​ place).
Answer #1

The equation is

Disk drives = 121.4 + 19.8 *Year

For Year 6,

Disk drives = 121.4 + 19.8*6 = 240.2

b)